OpenStudy (anonymous):

how do you solve x^2 - 4x?

OpenStudy (anonymous):

what means "solve" in this case? you do not have an equal sign

myininaya (myininaya):

solve?

OpenStudy (anonymous):

its # 4 im stuck on

myininaya (myininaya):

oh x^2-4x=0

myininaya (myininaya):

we can solve x^2-4x=0

myininaya (myininaya):

x(x-4)=0 x=0, x=4

OpenStudy (anonymous):

....... I feel stupid.

myininaya (myininaya):

oh you aren't stupid

OpenStudy (anonymous):

how about # 6?

myininaya (myininaya):

you corrected me earlier on cheat cheat lol

OpenStudy (anonymous):

OH Your the same person?! im new to this website!

myininaya (myininaya):

\[x^3-x^2-x=0\] \[x(x^2-x-1)=0\] \[x=0, x^2-x-1=0\]

myininaya (myininaya):

just use the quadratic formula for the second equation

OpenStudy (anonymous):

but it dont work

OpenStudy (anonymous):

why not?

myininaya (myininaya):

\[x=\frac{1 \pm \sqrt{1-4(1)(-1)}}{2(1)}=\frac{1 \pm \sqrt{5}}{2}\]

myininaya (myininaya):

so you have three real solutions!

OpenStudy (anonymous):

\[x=\frac{-b\pm\sqrt{b^2-4ac}}{2a}\] with \[a=1,b=-1,c=-1\]

OpenStudy (anonymous):

OH! i thought it had to be a whole #

OpenStudy (anonymous):

can u help me with # 10

myininaya (myininaya):

\[\frac{\frac{1}{x+1}}{x}=20\] \[\frac{1}{x(x+1)}=20\] \[1=20x(x+1)\] \[20(x^2+x)=1\] \[20x^2+20x-1=0\]

OpenStudy (anonymous):

put that into the quad. equation right?

myininaya (myininaya):

\[x=\frac{-20 \pm \sqrt{20^2-4(20)(-1)}}{20}=\frac{-20 \pm \sqrt{400+80}}{20}\] \[=\frac{-20 \pm \sqrt{480}}{20}=\frac{-20 \pm \sqrt{16 \cdot 30}}{20}=\frac{-20 \pm 4\sqrt{30}}{20}\] \[=\frac{4(-5 \pm \sqrt{30})}{4 \cdot 5}=\frac{5 \pm \sqrt{30}}{5}\]

myininaya (myininaya):

yes

OpenStudy (anonymous):

And # 14 kinda stumped me.

myininaya (myininaya):

\[2\sqrt{x}=x-3\] square both sides \[(2\sqrt{x})^2=(x-3)^2\] \[4x=x^2-6x+9\] \[x^2-6x-4x+9=0\] \[x^2-10x+9=0\] \[x^2-x-9x+9=0\] \[x(x-1)-9(x-1)=0\] \[(x-1)(x-9)=0=>x=1 or x=9\] ---- we must check these because one of them or both may not be a solution \[2\sqrt{1} \neq 1-3\] so x=1 is not a solution \[2\sqrt{9}=9-3\] so x=9 is a solution

OpenStudy (anonymous):

what about #15? that looked hard cause the cube root sign to me

myininaya (myininaya):

oops i already exited out of the attachment let me see if i can find it again

myininaya (myininaya):

\[\sqrt[3]{2x+1}-4=-1\] \[\sqrt[3]{2x+1}=4-1\] \[\sqrt[3]{2x+1}=3\] to gt rid of cube root cube both sides \[2x+1=3^3\] \[2x=27-1\] \[x=\frac{26}{2}=13\]

OpenStudy (anonymous):

and # 11?

myininaya (myininaya):

\[\frac{x}{x+2}-\frac{2}{2x-1}=\frac{1}{5}\] get rid of the fractions \[\frac{x}{x+2}5(x+2)(2x-1)-\frac{2}{2x+1}5(x+2)(2x-1)=\frac{1}{5}5(x+2)(2x-1)\] \[5x(2x-1)-10(x+2)=(x+2)(2x-1)\] \[10x^2-5x-10x-20=2x^2-x+4x-2\] \[10x^2-2x^2-5x+x-4x-10x-20+2=0\] \[8x^2-18x-18=0\] \[4x^2-9x-9=0\] use the quad formula

OpenStudy (anonymous):

# 15 seems so easy now. I feel dumb.

myininaya (myininaya):

i'm gonna take a break you should make a new thread if you have more questions

OpenStudy (anonymous):

good lord looka that latex!

OpenStudy (anonymous):

one more then i will stop. Please.

OpenStudy (anonymous):

you hit the question jackpot here myin

OpenStudy (anonymous):

and the next one is # 16.

OpenStudy (anonymous):

just post it. you will get lots of answers

myininaya (myininaya):

\[1+\frac{-2x}{\sqrt{25-x^2}}=0\] \[\sqrt{25-x^2}+(-2x)=0\] <--- i multiplied both sides by the sqrt(25-x^2} \[\sqrt{25-x^2}=2x\] square both sides \[25-x^2=(2x)^2\] \[25-x^2=4x^2\] \[4x^2+x^2-25=0\] \[5x^2-25=0\] \[x^2-5=0\] \[x^2=5=>x=\pm \sqrt{5}\]

myininaya (myininaya):

can i go now i need a break :(

OpenStudy (anonymous):

THANK YOU VERY MUCH. and yes you can leave. :D

myininaya (myininaya):

thanks for your permission lol

OpenStudy (anonymous):

LOL